By air, the distance from New York City to London is 2001 miles less than the distance from Los Angeles to Tokyo. If the total of these two distances is 8939 miles, find the distance from Los Angeles to Tokyo.
I would appreciate any help!! I'm really stuck on this one!
Thank you!

Answer by stanbon(75887)

By air, the distance from New York City to London is 2001 miles less than the distance from Los Angeles to Tokyo. If the total of these two distances is 8939 miles, find the distance from Los Angeles to Tokyo.
-----
x = y-2001
x+y = 8939
----
Rearrange:
x-y = -2001
x+y = 8939
---
Add and solve for "x":
2x = 7938
x = 3969 (distance NY to London)
---
y = 8939-3969 = 4979 (distance LA to Tokyo)
==============
Cheers,
Stan H.